OpenStudy (anonymous):

square root of x + 2 equals 10 - x

OpenStudy (anonymous):

umm....

OpenStudy (anonymous):

I think this is what you need, hope I'm right ! root x + 2 = 10-x root x = 8-x x= (8-x)^2 x= 64-16 x +x^2 x^2 -17x +64 = 0

OpenStudy (anonymous):

x=7